OpenStudy (anonymous):

Trigonometry: tan(theta/2 - pi/6) = 1 I'm a bit confused with how to start off on this one? considering theta is there and I don't know its value...

OpenStudy (anonymous):

I am supposed to solve for theta, but ya know lol. Difference formula doesn't seem to apply

OpenStudy (anonymous):

We have : \[\tan\alpha=1\iff \alpha=\frac\pi4+2k\pi\]

OpenStudy (anonymous):

Ooooh that seems familiar.

OpenStudy (anonymous):

theta/2 - pi/6 = pi/4 +2kpi ?

OpenStudy (anonymous):

I solve that?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

\[\theta-\frac\pi6=\frac\pi4+2k\pi\] So : \[\theta=\frac\pi4+\frac\pi6+2k\pi\]

OpenStudy (anonymous):

theta = 5pi/12 + 4kpi?

OpenStudy (anonymous):

Wait, how did you get rid of the denominator, 2, of theta?

OpenStudy (anonymous):

It is : \[\frac\theta2\] I didn't notice that !

OpenStudy (anonymous):

Oh it is xD No problem!

OpenStudy (anonymous):

I just don't really use the equation thing, it's a bit frustrating to me.

OpenStudy (anonymous):

Wait, and another thing.. Isn't the period of tangent pi? not 2pi? Or am I mistaken?

OpenStudy (anonymous):

You are right ! The period is pi !

OpenStudy (anonymous):

Alright, well thanks for the help!